Jim purchased 2 pieces of property that had been foreclosed on by a local bank for a total of $330,000. When he sold these properties, on the first piece Jim earned a profit of 12% but on the second piece he lost 8%. his total profit was $18,000. How much did Jim pay for each property?
Answer by stanbon(75887)

Jim purchased 2 pieces of property that had been foreclosed on by a local bank for a total of $330,000. When he sold these properties, on the first piece Jim earned a profit of 12% but on the second piece he lost 8%. his total profit was $18,000. How much did Jim pay for each property?
-----
Equation:
Interest - Interest = Interest
---------
0.12x - 0.08(330,000-x) = 18000
--------
12x - 8*330000+8x = 1800000
---------------------
20x = 1800000+8*330000
-------
20x = 4440000
---
x = $222,000 (amt invested at 12%
---
333,000- 222,000 = $111,000 (amt invested at 8%)
------
Cheers,
Stan H.
-------